Ashley went to the gas station on Saturday and bought 2 Gatorades as well as 2 Red bulls and her total cost was $ 10. she returned the next saturday and bought 3 gatorades and 4 red bulls and this time her total cost was $18. how much did a gatorade n red bull cost?
1
Expert's answer
2012-08-16T09:13:04-0400
Let one gatorade costs x and one red bull y
We can construct the following two equations:
2x+2y=10& => x=5-y
3x+4y=18

3(5-y)+4y=18
15-3y+4y=18
y=3
Therefore x=5-3=2

Answer: $2, $3